# Larger Is Not Always Better: Exploring Small Open-source Language Models in Logging Statement Generation

### v2

@@ 5. Discussion
Based on previous experiments, this section analyzes the strengths of using SOLMs and elicits potential future research directions to advance logging practice.
@@ 5.1.1. Efficiency and Cost-Effectiveness
SOLMs exhibit remarkable efficiency and cost-effectiveness in automated logging, making them a compelling alternative to LLMs. Unlike LLMs, which often require extensive computational resources, including thousands of GPU hours for training and inference, SOLMs achieve comparable performance with significantly lower resource demands. For instance, our experiments demonstrate that the Qwen2.5-coder-14B model, with 14 billion parameters, can be fine-tuned within 6 hours using a single Nvidia A100 GPU, producing high-quality logging statements with a PA of 66.20% and a ROUGE-L score of 46.51% (Table 5). This efficiency reduces hardware costs and energy consumption, aligning with sustainable computing goals.
@@ 6.1. Internal Validity
Selection of hyperparameters for fine-tuning. A potential threat to internal validity lies in the selection of hyperparameters for fine-tuning the SOLMs (e.g., learning rate, batch size, prefix length for prefix tuning, rank for LoRA). The study utilized recommended hyperparameters from official sources due to their proven effectiveness. However, these hyperparameters may not be optimal for all models or datasets, potentially introducing bias in the performance results. Suboptimal hyperparameter choices could lead to underperformance or overfitting, affecting the observed effectiveness of SOLMs compared to baseline LLMs or existing methods. To mitigate this, we conducted preliminary experiments to validate the chosen hyperparameters on a subset of the AL-Bench dataset, ensuring reasonable performance. Nonetheless, a more exhaustive hyperparameter search may further enhance model performance and reduce this threat. The potential data leakage. A potential threat to the internal validity of our study is the possibility that the AL-Bench dataset, which comprises logging statements from 10 widely used open-source projects, may have been included in the pre-training corpora of the SOLMs or LLMs evaluated. Since these models are pre-trained on large-scale datasets, often including publicly available code repositories from platforms like GitHub, there is a risk that some or all of the AL-Bench projects were part of their training data. Such data leakage could artificially inflate the performance of these models, particularly in zero-shot or few-shot settings, potentially skewing our results and conclusions. To mitigate this threat, we carefully analyzed the performance of base (non-fine-tuned) models in our experiments. The results, as shown in Table 3, indicate that base models, such as Qwen2.5-coder (PA: 4.30%), exhibit significantly lower performance compared to their fine-tuned counterparts (PA: 62.40%). This substantial performance gap suggests that the base models, despite potential exposure to AL-Bench data during pre-training, do not inherently possess the task-specific knowledge required for high-quality automated logging. Instead, the superior performance of fine-tuned SOLMs is primarily attributable to our fine-tuning strategies (e.g., LoRA and RAG), which adapt the models to the specific logging task and dataset. Thus, we argue that any potential data leakage has minimal impact on the validity of our conclusions, as the observed improvements stem from task-specific fine-tuning rather than pre-existing knowledge from pre-training. Nonetheless, to further address this threat in future work, we recommend evaluating models on proprietary or newly created datasets that are guaranteed to be absent from pre-training corpora.
@@ 6.2. External Validity
The representativeness of the dataset. A potential threat to generalizability is that the AL-Bench dataset, comprising 10 open-source projects, may not fully represent logging practices in proprietary codebases. We mitigated this by selecting projects from diverse domains (e.g., task scheduling, messaging systems, IoT platforms), ensuring broad coverage of logging requirements. However, similar to prior work (Li et al., 2024a), our dataset is predominantly Java-based, which may limit the generalizability of SOLM’s performance in generating log statements for other programming languages. This language-specific focus could restrict insights into how SOLM performs across diverse language ecosystems, potentially affecting its applicability in non-Java contexts. The selection of SOLM. Another potential threat to the generalizability of our findings lies in the selection of specific SOLMs evaluated in this study. While these models were chosen based on their established performance in software engineering tasks, they may not fully represent the diversity of available SOLMs or future advancements in model architectures. For instance, other SOLMs with different pre-training datasets, architectural designs (e.g., transformer variants or mixture-of-experts models), or domain-specific optimizations might exhibit varying performance in automated logging. To mitigate this threat, we selected models with broad applicability in code-related tasks and ensured they were fine-tuned using techniques (e.g., LoRA and RAG) to align with logging-specific requirements. However, future work should explore a wider range of SOLMs, including those with different training corpora or specialized architectures, to validate the robustness of our findings across diverse model ecosystems.
@@ 6.3. Construct Validity
Adequacy of evaluation metrics for logging quality. A potential threat to construct validity lies in whether the chosen evaluation metrics fully capture the quality of generated logging statements. These metrics primarily assess syntactic similarity to ground-truth logs (e.g., ROUGE-L for text similarity and correctness of logging location (PA). However, high-quality logging statements must also provide actionable insights for developers, such as facilitating debugging or system monitoring, which may not be fully reflected by these metrics. For instance, a generated log might score high on ROUGE-L due to textual similarity but fail to capture critical runtime context (e.g., omitting key variables or using an inappropriate verbosity level). To mitigate this, we incorporated the LLM-as-a-judge approach to assess overall quality holistically. Nevertheless, future work could include developer-centric evaluations, such as user studies, to validate the practical utility of generated logs. Representativeness of the LLM judger result. The use of an LLM to evaluate the quality of generated logging statements introduces a construct validity threat if the LLM’s judgments do not align with human developer preferences. While the LLM-as-a-judge approach has shown promise in software engineering tasks (Wang et al., 2025), its scoring may not fully capture nuanced aspects of log quality, such as clarity, relevance to specific debugging scenarios, or adherence to project-specific logging conventions. Misalignment between LLM and human judgments could lead to over- or underestimation of SOLM performance. Therefore, incorporating human evaluations or domain-specific rubrics in future work could enhance the alignment between the LLM judger and practical logging needs.
@@ 7.1. Automated Logging
Traditionally, the automation of logging statements is divided into two primary stages (Chen and Jiang, 2021; He et al., 2021): the identification of logging locations and the creation of logging statements. These stages are respectively denoted as where to log and what to log (Zhong et al., 2025). In addressing the complexities associated with determining where to log, various methodologies have been investigated by researchers to identify appropriate logging locations within source code (Jia et al., 2018; Li et al., 2020a; Yao et al., 2018; Zhao et al., 2017; Zhu et al., 2015; Yuan et al., 2012b; Li et al., 2018). Regarding what to log, the generation of logging statements is usually segmented into three specific subtasks: the generation of logging text (Ding et al., 2022), the selection of logging variables (Liu et al., 2019; Yuan et al., 2012c), and the prediction of the logging level (Li et al., 2017, 2021a; Liu et al., 2022; Mizouchi et al., 2019).

### v3

@@ 5. Discussion
5.1. In-depth Analysis: Case Study of Success and Failure
5.1.1. Successful Case
Figure 7. A Case of Generated Logging Statement from Multiple Models.
Figure 7 presents a case study comparing logging statements generated by different models for an error logging scenario in a Java method handling a DuplicateKeyException. The ground truth logging statement is log.error("Update alert group error, groupName:{}"), alertGroup.getGroupName(), ex). Among the LLMs, Claude3.7sonnet-RAG generates a statement that closely resembles the ground truth but includes an additional variable (id), which may add unnecessary verbosity. Deepseek-coder-v3-RAG and GPT4o-RAG produce statements with missing or misused variables (e.g., using groupName directly instead of alertGroup.getGroupName()), reducing their contextual accuracy. LLaMA3.1-405B-RAG omits the exception variable (ex), limiting its diagnostic utility. In contrast, fine-tuned SOLMs demonstrate superior performance: Mistral-7B-RAG-LoRA and CodeLlama-13B-RAG-LoRA generate statements identical to the ground truth, ensuring both accuracy and relevance. Qwen2.5-coder-14B-RAG-LoRA adds minor additional text (”already exist”) but retains all critical components, closely aligning with the ground truth.
5.1.2. Failure Case
Figure 8. A Case Study on the Failure to Invoke Out-of-Context Methods in Log Generation.
Figure 8 presents a failure case study comparing logging statements generated by different models. This case study reveals a critical limitation of the models: an inability to generate logging statements that invoke methods not explicitly visible within the immediate code snippet. The diagnostic purpose of the ground-truth log is to capture the output of the .explainQuery() method, which is chained to the dorisAssert.query(query) call. However, most models failed to reason about or infer the existence of this critical but locally unseen method. Instead, they adopted safer, more conservative strategies based only on visible artifacts. For instance, several models correctly identified the logging location but defaulted to logging the simple query variable, demonstrating a failure to explore the object’s potential API. More severely, models like Qwen2.5-coder abandoned contextual analysis altogether, generating a generic and uninformative ‘Success!’ message. This tendency to avoid invoking out-of-context methods severely limits their ability to create truly insightful, diagnostic logs and underscores the necessity of fine-tuning on domain-specific API usages to overcome this limitation. Future work could address this limitation by enriching the model’s input with structural code representations (e.g., call graphs from static analysis) or by designing more sophisticated, type-aware retrieval mechanisms for RAG that can fetch relevant API usage patterns on demand.
Figure 9. A Case Study on Distraction by Normal Execution Flow in Complex Methods during Log Generation.
Figure 9 illustrates a critical failure mode in navigating long, complex methods with high logging density. The ground-truth log is situated within a catch (IOException e) block, making its purpose unequivocally tied to error handling. However, the method’s length and the presence of five other log statements appear to have primed most models for a different task. Distracted by this high density of existing logs within the try block, models like Qwen2.5-coder seem to have lowered their activation threshold for what constitutes a logging event, becoming overly willing to log the incremental progress of the loop. Consequently, they ignored the distinct exceptional path and prematurely generated logs related to the loop’s normal termination, with messages such as ‘Reached end of block iteration’. This failure likely points to a fundamental limitation of attention-based models, which excel at identifying patterns within the provided context but struggle with reasoning about or ‘imagining’ external APIs not explicitly mentioned. It suggests that without specialized fine-tuning on a project’s entire API surface, SOLMs may default to a ‘what you see is what you get’ behavior, limiting their ability to generate truly insightful, diagnostic code.
The failure to prioritize the exceptional path is further highlighted by the contrasting success of models like Claude3.7, which correctly identified the catch block despite the noisy, log-dense environment of the main logic. The inability of most models to navigate this complexity represents a severe limitation for a practical automated logging tool, as a primary function of logging is to capture unexpected errors. This case suggests that models can be desensitized by existing logs, leading them to misjudge the relative importance of the normal paths versus critical error-handling blocks. Future work should therefore focus on improving models’ structural awareness and their ability to reason about control flow priority in lengthy, real-world methods that already contain logging. This case highlights a potential ‘attention bias’ in log-dense code. The model, when exposed to numerous logging examples for normal execution paths, appears to lower its threshold for what constitutes a log-worthy event. It consequently fails to assign a higher priority to the semantically critical but less frequent exceptional path in the catch block. This suggests that future research should focus not just on code semantics, but on teaching models to reason about control flow priority and the relative importance of different execution paths.
5.2. Implications and Advice
Our empirical analysis showcases the potential of fine-tuned SOLMs. Based on our findings, we identify three key implications and provide actionable advice to guide future research and practice in automated logging statement generation.
Adopt a Synergistic Optimization Blueprint. Our results consistently demonstrate that the optimal performance of SOLMs is not achieved through any single technique but through a deliberate, multi-stage optimization process. We found that while ‘instruct-tuned’ models provide a superior foundation over ‘base’ models, their true potential is only unlocked via task-specific fine-tuning, for which LoRA emerged as the most effective and stable PEFT technique. Furthermore, the performance of a LoRA-tuned model is maximally amplified when coupled with a RAG pipeline, which provides critical instance-specific context at inference time. The synergy between task-level specialization (from LoRA) and instance-level specialization (from RAG) consistently yields the best results. This suggests a clear blueprint for developing high-performance, customized code generation tools.
Implication 1. Future research and development of SOLM-based tools should adopt a holistic optimization strategy.
Re-evaluate the Performance-vs-Resource Paradigm. A central finding of our work challenges the prevailing ”larger is better” paradigm. We demonstrate that a well-optimized, sub-14B parameter SOLM can significantly outperform massive, proprietary LLMs and existing SOTA tools in both logging location and statement quality. This indicates that for specialized SE tasks, targeted adaptation is more critical than raw model scale. However, this does not imply that the smallest models are sufficient; our findings reveal a performance ‘sweet spot’, with models above 3B parameters showing more consistent and meaningful performance gains. This trade-off is further nuanced in complex scenarios; on long code snippets, fine-tuned SOLMs show a greater decline in positioning accuracy than LLMs but achieve substantially larger improvements in content quality, proving more adept at leveraging rich context.
Implication 2. Practitioners should shift from defaulting to the largest available models and instead identify a cost-effective ‘sweet spot’ for SOLM size.
Prioritize Data Strategy and User Experience for Robustness. Our study instills confidence in the real-world deployability of SOLMs, which exhibit strong generalization to unseen repositories. However, we found that this robustness is not inherent but is significantly influenced by data and context. The generalization capability of a model is greatly enhanced when it is fine-tuned on a corpus with consistent logging standards, such as those found across Apache Foundation projects. Furthermore, practitioners must be cognizant of the performance trade-offs introduced by code length, where positioning accuracy degrades while content quality improves. The stability of the LoRA+RAG strategy across these varied conditions underscores its reliability.
Implication 3. To build robust and generalizable tools, practitioners should prioritize the curation of fine-tuning datasets that reflect consistent coding standards.
@@ 5.3.1. Efficiency and Cost-Effectiveness
From a practical enterprise standpoint, SOLMs present a highly efficient and cost-effective solution for automated logging compared to strategies centered on LLMs. An organization seeking a customized logging tool faces a trade-off: using a commercial LLM via its API involves continuous, per-token operational costs that can become substantial at scale, alongside significant privacy risks from sending proprietary code to a third party. Alternatively, self-hosting and fine-tuning a massive open-source LLM requires a prohibitive investment in computational resources. Our approach offers a compelling middle ground. As demonstrated in our experiments, an SOLM like Qwen2.5-coder-14B can be fine-tuned on enterprise-specific data in under 6 hours on a single A100 GPU to achieve state-of-the-art results (Table 5). This makes the initial customization cost manageable and, more importantly, enables local deployment for inference. Such deployment drastically lowers the long-term cost and energy consumption while completely preserving data privacy, offering a practical and sustainable path for adopting automated logging tools.
@@ 5.4.2. Addressing Dynamic Logging Requirements
5.4.3. Enhancing Long-Context Reasoning for Accurate Positioning
Our empirical results revealed a key trade-off: while fine-tuned SOLMs excel at leveraging the richer context of long code to improve logging content quality, their accuracy in determining the precise logging location degrades more significantly compared to larger LLMs. This highlights an important direction for future research in enhancing the long-context reasoning of SOLMs for precise automated logging tasks. Future work could explore architectural enhancements that integrate explicit code structure representations like Code Property Graphs (CPGs). Successfully addressing this challenge would significantly improve the robustness of SOLMs, further solidifying their role as a scalable and reliable solution for automated logging in large-scale industrial projects where complex methods are common.
@@ 6.1. Internal Validity
6.1.1. Selection of hyperparameters for fine-tuning.
A potential threat to internal validity lies in the selection of hyperparameters for fine-tuning the SOLMs (e.g., learning rate, batch size, prefix length for prefix tuning, rank for LoRA). The study utilized recommended hyperparameters from official sources due to their proven effectiveness. However, these hyperparameters may not be optimal for all models or datasets, potentially introducing bias in the performance results. Suboptimal hyperparameter choices could lead to underperformance or overfitting, affecting the observed effectiveness of SOLMs compared to baseline LLMs or existing methods. To mitigate this, we conducted preliminary experiments to validate the chosen hyperparameters on a subset of the AL-Bench dataset, ensuring reasonable performance. Nonetheless, a more exhaustive hyperparameter search may further enhance model performance and reduce this threat.
6.1.2. The potential data leakage.
A significant threat to the internal validity of our study is the possibility that the AL-Bench dataset may have been part of the large-scale corpora used to pre-train the evaluated SOLMs and LLMs. Since AL-Bench consists of popular open-source projects, it is plausible that its code was included in web scrapes from platforms like GitHub. If such data leakage occurred, the models’ performance might be artificially inflated, stemming from the recall of memorized examples rather than from genuine learning of the automated logging task. We acknowledge this critical issue and present several lines of evidence from our experiments that mitigate this threat.
First, we observed a substantial performance gap between non-fine-tuned base models and their fine-tuned counterparts (e.g., Qwen2.5-coder’s Position Accuracy surged from 4.30% to 62.40% after fine-tuning). This demonstrates that our task-specific fine-tuning, not pre-existing knowledge, is the primary driver of performance. Second, our cross-repository generalization experiment (RQ4.1) showed that models trained on one set of projects perform robustly on entirely new, unseen ones. This ability to generalize to new codebases is inconsistent with a simple RAG-based project-specific memorization hypothesis. Finally, our analysis on long code snippets (RQ4.2) revealed a nuanced trade-off: richer context made log position more difficult but improved log content quality. This complex behavior suggests a genuine reasoning process where models leverage semantic context, rather than a simple recall mechanism, which would likely have improved all metrics uniformly.
Collectively, this evidence strongly suggests that our findings are robust and that the observed performance stems from effective, task-specific adaptation rather than data leakage. Nonetheless, to further address this threat in future work, we recommend evaluating models on proprietary or newly created datasets that are guaranteed to be absent from pre-training corpora.
@@ 6.2. External Validity
6.2.1. The representativeness of the dataset.
A potential threat to generalizability is that the AL-Bench dataset, comprising 10 open-source projects, may not fully represent logging practices in proprietary codebases. We mitigated this by selecting projects from diverse domains (e.g., task scheduling, messaging systems, IoT platforms), ensuring broad coverage of logging requirements. However, similar to prior work (Li et al., 2024a), our study focuses exclusively on Java-based projects, which may limit the generalizability of our findings on SOLM performance to other programming languages. This language-specific focus could restrict insights into how SOLMs perform across diverse language ecosystems, potentially affecting their applicability in non-Java contexts.
6.2.2. The selection of SOLM.
Another potential threat to the generalizability of our findings lies in the selection of specific SOLMs evaluated in this study. While these models were chosen based on their established performance in software engineering tasks, they may not fully represent the diversity of available SOLMs or future advancements in model architectures. For instance, other SOLMs with different pre-training datasets, architectural designs (e.g., transformer variants or mixture-of-experts models), or domain-specific optimizations might exhibit varying performance in automated logging. To mitigate this threat, we selected models with broad applicability in code-related tasks and ensured they were fine-tuned using techniques (e.g., LoRA and RAG) to align with logging-specific requirements. However, future work should explore a wider range of SOLMs, including those with different training corpora or specialized architectures, to validate the robustness of our findings across diverse model ecosystems.
@@ 6.3. Construct Validity
6.3.1. Adequacy of evaluation metrics for logging quality.
A potential threat to construct validity lies in whether the chosen evaluation metrics fully capture the quality of generated logging statements. We identified two key limitations in the traditional metrics used. First, metrics for static text evaluation, such as BLEU-4 and ROUGE-L, are confined to lexical similarity and often overlook semantic consistency. A generated statement might be semantically identical or superior to the ground truth but use different wording, causing it to receive a low score unfairly. Second, the PA metric requires an exact match with the ground-truth location, which is overly rigid. In practice, multiple locations within a code block might be functionally equivalent for logging, but the PA metric would penalize these valid, alternative placements. To mitigate these inherent metric-based limitations, we incorporated the LLM-as-a-judge approach to holistically assess the overall quality of the generated logging statements. Unlike rigid metrics, the LLM judges evaluate each statement based on a comprehensive scoring guideline that prioritizes semantic relevance, contextual appropriateness, and syntactic correctness. This method allows for a more nuanced assessment; for instance, a statement that is semantically correct but lexically different, or one placed in a functionally equivalent but not identical location, may still receive a high score. This qualitative evaluation, therefore, compensates for the weaknesses of traditional metrics and provides a more practical and robust assessment of a model’s true performance.
6.3.2. Representativeness of the LLM judger result.
The use of an LLM to evaluate the quality of generated logging statements introduces a construct validity threat if the LLM’s judgments do not align with human developer preferences. While the LLM-as-a-judge approach has shown promise in software engineering tasks (Wang et al., 2025), its scoring may not fully capture nuanced aspects of log quality, such as clarity, relevance to specific debugging scenarios, or adherence to project-specific logging conventions. Misalignment between LLM and human judgments could lead to over- or underestimation of SOLM performance. Therefore, incorporating human evaluations or domain-specific rubrics in future work could enhance the alignment between the LLM judger and practical logging needs.
@@ 7.1. Automated Logging
Traditionally, the automation of logging statements is divided into two primary stages (Chen and Jiang, 2021; He et al., 2021): the identification of logging locations and the creation of logging statements. These stages are respectively denoted as where to log and what to log (Zhong et al., 2025). In addressing the complexities associated with determining where to log, various methodologies have been investigated by researchers to identify appropriate logging locations within source code (Jia et al., 2018; Li et al., 2020; Yao et al., 2018; Zhao et al., 2017; Zhu et al., 2015; Yuan et al., 2012b; Li et al., 2018). Regarding what to log, the generation of logging statements is usually segmented into three specific subtasks: the generation of logging text (Ding et al., 2022), the selection of logging variables (Liu et al., 2019; Yuan et al., 2012c), and the prediction of the logging level (Li et al., 2017, 2021a; Liu et al., 2022; Mizouchi et al., 2019).